One-fourth of a number is 3 greater than one-sixth of it. Find the number.
Answer by ptaylor(2198)

Let x=the number
x/4=one-fourth of the number
x/6=one-sixth of the number
Now we are told the following:
x/4=(x/6)+3 multiply each term by 12 (LCD)
3x=2x+36 subtract 2x from each side
3x-2x=2x-2x+36 collect like terms
x=36-----------------------------------the number
CK
(1/4)*36=(1/6)*36+3
9=6+3
9=9
Does this help??------------------ptaylor